OpenStudy (anonymous):

Solve 4(-2)^5+6(-2)^3+(-2)^2-10(-2)+5 Show work please. Thank you!

OpenStudy (anonymous):

= \[4(-32) + 6(-8) +4 +20+5 \] = -128-48+4+20+5 = -147

OpenStudy (anonymous):

Thanks! ^_^

OpenStudy (anonymous):

Bianca Wesbrook :)

OpenStudy (anonymous):

would u rather add me on urs

OpenStudy (anonymous):

It might be hard to find you, so idk. I will try to find you though. :D

OpenStudy (anonymous):

my profile pic is the same on here

OpenStudy (anonymous):

okay cool! ^_^ shouldnt be that hard then. I'll find you on fb. btw, wat is your profile pic? Kinda small on this site...a mission trip?

OpenStudy (anonymous):

it is a egine just dont know wat kind!!(P(:

OpenStudy (anonymous):

okay, i really wasnt sure ! Looks like ur doing something on a mission trip, sorry! :P lol, anyway, i'll remember to find you on fb this weekend. I have to go now, i have hw.

OpenStudy (anonymous):

try matthew porter on fb

OpenStudy (anonymous):

okay i will this weekend. :)

OpenStudy (anonymous):

kay bye bianca(::P